OpenStudy (anonymous):

Word Problem: Application of Quadratic Equations - A rectangle shaped x-ray film has an area of 80 square inches. The length is 2 inches longer then the width. Find the width and length.

OpenStudy (anonymous):

let width = x, length = x+2. Then x(x+2)=80, so x^2 +2x - 80 = 0

OpenStudy (anonymous):

\[A = lw\]\[l = w+2\] \[A = (w)(w+2) = 80\]\[w ^{2}+2w = 80\]\[w^{2}+2w-80=0\] Quadratic Formula if you can't figure it out.

OpenStudy (anonymous):

a = 1, b=2, c = -80 into the quadratic formula

OpenStudy (anonymous):

You end up with 8 and -10. It has to be positive so \[w=8\]\[l=10\]

OpenStudy (anonymous):

Ohhhhhh thank you... I am seriously struggling with these words problems!